Find the volume of the solid generated by revolving the region bounded by the graphs of the equations about the indicated line.

y = 2 − x^2 ,y = 0; the line x = 2.
Answer #1

y=2-x^2 1- x=2 y=0 5 (1.414, 0) (2,0) - 1 (2, -2) --2 --3 --4 >>Given, y=2-002 y=0 and 2 = 2 The bounded region by above equations is given is the graph. This region when ratated about z-ax